OpenStudy (anonymous):

The triangles shown are similar. What is the value of x? A. 9 B. 15 C. 12 D. 10

OpenStudy (anonymous):

http://prntscr.com/6utj1f

OpenStudy (anonymous):

@Godlovesme

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

@CrystalBurley

OpenStudy (dmndlife24):

You can set up a proportion..like this If the leg of the little triangle is 3 units, the corresponding triangle leg will be proportional, meaning its length can double or triple. In this case the corresponding triangles leg is 6 units...so the length doubled

OpenStudy (butterflydreamer):

your triangles are similar :) Given the diagram that you have, since they've given us 3 sides. This means that ALL three sides must be in proportion. So 6/3 = 8/4 = x/5 Therefore what would x be given that the ratios must always be the same :) ?

OpenStudy (anonymous):

can i c the triangles first?

OpenStudy (dmndlife24):

3/6 = 4/8 = 5/x simplify... 1/2 = 5/x Solve for x

OpenStudy (anonymous):

First of all how in the world can u solve the awnser without seeing the triangle it can b any kind of triangle so how in the world can u solve it

OpenStudy (anonymous):

http://prntscr.com/6utj1f

OpenStudy (dmndlife24):

http://prntscr.com/6utpax

OpenStudy (anonymous):

^^Thx 2 both of u ^^

OpenStudy (anonymous):

I am still kinda confused with the asnwer

OpenStudy (dmndlife24):

To find x, we have to set up a proportion

OpenStudy (butterflydreamer):

basically what does 6/3 = ? 8/4 = ? They both equal to 2 right? So since the proportion/ratio is 2 : 1 (basically the bigger triangle, has sides which are DOUBLE in length) . Then to find x/5, we know x has to be double the smaller side! The smaller side is from the SMALLER triangle (so the side with length 5) so to find x, we DOUBLE 5 :) So x = 2*5 = ?

OpenStudy (anonymous):

10

OpenStudy (dmndlife24):

Yes

OpenStudy (anonymous):

1) Since we know 2 sides of this triangle, we will use the pythagorean theorem to solve for x. 2)Substitue the two known sides into the pythagorean theorem's formula: A² + B² = C² 8² + 6² = x² 100 = x²

OpenStudy (anonymous):

=10 sorry

OpenStudy (dmndlife24):

So basically what they mean when they say "similar" is that all corresponding sides are in proportion. This means that one could be larger than the other or smaller, but also similar. In this case, side 3 corresponds to 6 and 4 to 8 and 5 to x 3/6 = 4/8 = 5/x 1/2 = 5/x (5 over what equals 1/2?) x = 10

OpenStudy (anonymous):

^_^Yes^_^

OpenStudy (dmndlife24):

@TheAsker2002 do you understand?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thanks guys idk who to give medal to though

OpenStudy (anonymous):

dmndlife24 deserves it or we could just all give every1 medal

OpenStudy (anonymous):

I have more i need help with hold on

OpenStudy (anonymous):

iam here 2 help bff